OpenStudy (anonymous):

Let theta be an angle in quadrant II such that cot(theta)=-4/3 Find the exact values of sin and sec

OpenStudy (anonymous):



OpenStudy (anonymous):

the only thing you have to pay attention to is what quadrant you are in

OpenStudy (anonymous):

since you are in quadrant II, you know that cosine is negative and sine is positive. then you can read off the answers from the triangle

OpenStudy (anonymous):

easy 3 - 4 - 5 right triangle in this case, so \[\sin(\theta)=\frac{3}{5}\] etc

OpenStudy (anonymous):

maybe this would be a better picture